# MixAT: Combining Continuous and Discrete Adversarial Training for LLMs

### v1

@@ B.2 Discussion on Randomness
Randomness of attack generation All of the adversarial attacks we use in our experiments have a random component in their generation and evaluation process, since including random sampling from a distribution (GCG) or querying LLMs with a non-zero temperature (TAP, PAIR, AutoDAN). The adversarial attacks are evaluated by LLM-as-a-judge [31], which might cause inaccuracies. Randomness can yield different results when evaluating the robustness of the same model. To measure the extent of this randomness, we generate the adversarial samples for a single model multiple times and evaluate the robustness of the model on each set of samples. The results in Table 4 show that the variance in the ASR scores is relatively low in the case of Zephyr-7B Base and MixAT model, while CAT results (PAP, TAP, PAIR) have significant randomness. This can stem from that the evaluation results of CAT indicating moderate robustness, which can only defend against weaker adversarial attacks. Therefore, this huge variance might indicate one limitation of LLM adversarial attacks, namely that it is difficult to generate similarly strong adversarial examples.
@@
R1 64.4 97.5 54.5 88.4 2.5 10.0 12.5
R2 63.8 97.5 53.4 74.0 0.0 12.5 12.5
R3 64.2 95.0 53.0 62.8 7.5 7.5 12.5
Avg 64.1 96.7 53.6 75.1 3.3 10.0 12.5
Std 0.31 1.44 0.79 12.83 3.82 2.50 0.00
Randomness of training We also examine the effect of randomness in the training process. We train the same model with different random seeds and evaluate the robustness of the models against adversarial attacks. The results in Table 5 show that the variance induced by the randomness in the MixAT training procedure of Zephyr-7B has a relatively small magnitude, except the XSTest results. We would like to examine the causes of XSTest randomness in future work.
@@ B.3 Impact of Model Quantization
To examine the impact of model quantization, we train several models without 4-bit quantization and evaluate both the base and 4-bit quantized versions of them. The results (Table 6) suggest, that models trained without 4-bit quantization are less robust than the models which were trained in 4-bit quantized form (higher ALO-ASR).
@@
Model Quantization Utility Benchmarks [%] Attack Success Rate [%]
@@
Zephyr-7B Base Eval 81.0 55.2 56.2 100.0 61.4 98.8 85.0 87.5 85.0 97.5 90.0 85.0 100.0 100.0
None 81.3 57.3 58.1 100.0 61.3 98.8 77.5 72.5 90.0 100.0 95.0 90.0 100.0 100.0
Zephyr-7B CAT Train + Eval 78.2 50.5 54.5 95.0 53.7 50.0 0.0 25.0 27.5 55.0 0.0 12.5 0.0 67.5
Eval 77.7 50.4 55.0 90.0 54.2 47.6 0.0 45.0 55.0 65.0 0.0 12.5 0.0 87.5
None 79.2 51.5 56.9 90.0 55.7 50.8 0.0 52.5 52.5 75.0 0.0 25.0 2.5 90.0
Zephyr-7B MixAT Train + Eval 81.4 54.0 55.8 97.5 54.3 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
Eval 81.9 54.4 56.3 92.5 53.3 55.2 0.0 2.5 2.5 12.5 0.0 5.0 0.0 15.0
None 82.4 57.4 57.9 92.5 53.7 54.8 0.0 10.0 10.0 12.5 0.0 5.0 0.0 22.5
@@ B.4 Evaluating on more samples
In all our experiments we evaluate the models on a subset of 40 samples from the HarmBench test set. To assess the generalization of our models, we evaluate some of the Zephyr-7B models on the full test set of the HarmBench dataset. However, we still exclude the copyright samples as before, and we have a total of 240 samples. Since generating GCG attacks for all 240 samples would take 6 days on a single A100 GPU for each evaluated model, we only evaluate the models using the other attacks. The results in Table 7 show that the models achieve similar robustness scores on the full test set as on the subset used in the main experiments. This indicates that evaluating on the 40-sample subset is a good proxy for the full HarmBench test set.
We also evaluate the robustness of the Zephyr models on a subset of 40 samples from the harmful prompts of the XSTest dataset. The results in Table 8 show that the general robustness trends are similar across the two datasets, when considering the same attack.
@@
Model Harmbench Samples Attack Success Rate [%]
@@ B.5 Transferability of attacks
Prior work [4] showed that adversarial attacks generated for non-adversarially-trained models transfer well to many other models. Here, we want to assess if adversarial attacks generated for adversarially-trained models exhibit the same trend. To do this, we generate adversarial samples using TAP, PAIR, and GCG attacks for different Zephyr-based models — undefended Z-Base, as well as, Z-CAT, and our Z-MixAT. We evaluate these sets of generated attacks on a large set of models both based on Zephyr and Llama3 and compare their efficiency versus Direct Requests and attacks specifically generated for the models. The results in Table 9 show that almost all transferred attacks achieve worse performance compared to their targeted counterparts, with some transferred attacks being even less successful than the direct requests. Further, MixAT shows incredible resilience against all adversarial attacks, regardless of their origin.
@@
Defender Model Direct R. ASR [%] ASR [%] of TAP when target is ASR [%] of PAIR when target is ASR [%] of GCG when target is
@@ B.6 Utility benchmarks without chat template
Recent work highlights a failure mode in previous works, that evaluating LLMs on robustness and utility benchmark inconsistently with or without chat template might produce unrealistic results [7], since in real-world scenarios the settings of text generation do not depend on the content of prompts or whether they are malicious or ordinary requests. To enhance consistency, in this section we present ARC-E, ARC-C and MMLU results of LLama3 and Zephyr-7b with their default chat template Table 10.
@@
Evaluated Model Chat Utility Benchmarks [%]
@@ B.7 MixAT on other models
To demonstrate the generalizability of our method, MixAT has been evaluated on additional models. We aimed to assess MixAT on an adversarially less robust (Mistral-7B[33]) and more robust (Llama3.1-8B [25]) base models. The evaluation results are in Table 11. These are similar to our original findings based on Zephyr-7B, LLama3-8B and Qwen2.5 models. We observe that MixAT drastically increases the robustness of the base model in all of the cases, with a slight drop in utility.
@@
Model Utility Scores [%] $\uparrow$ Attack Success Rate [%] $\downarrow$
@@
Mistral-7B No Defense (HF) 79.7 49.9 52.5 100.0 60.0 99.2 80.0 77.5 90.0 95.0 95.0 80.0 100.0 100.0
CAT [7] (R) 79.8 50.0 52.8 90.0 54.9 68.8 2.5 60.0 80.0 80.0 7.5 37.5 32.5 95.0
PAP-AT 80.1 51.3 52.9 97.5 52.8 85.2 12.5 10.0 50.0 65.0 30.0 50.0 100.0 100.0
MixAT 79.8 50.3 52.8 92.5 52.1 55.2 2.5 7.5 22.5 25.0 0.0 40.0 10.0 52.5
Llama3.1-8B No Defense (HF) 80.2 49.9 64.0 100.0 75.3 94.4 30.0 67.5 55.0 77.5 97.5 57.5 100.0 100.0
CAT [7] (R) 80.6 51.7 61.5 90.0 68.1 73.2 0.0 37.5 50.0 75.0 2.5 20.0 5.0 82.5
PAP-AT 81.3 51.1 64.1 100.0 65.3 92.4 15.0 5.0 12.5 22.5 7.5 55.0 35.0 67.5
MixAT 81.3 51.5 63.4 92.5 62.4 59.2 0.0 0.0 2.5 10.0 0.0 5.0 0.0 12.5
@@ B.8 MixAT Static training ablation results
We present the results of opur ablation using statically generated adversarial attacks in the training procedure of MixAT. The evaluation results are in Table 12. We observe that while the static model has slightly higher utility, this comes at the cost of much worse robustness. This reaffirms the effectiveness of MixAT’s dynamic attack component.
@@
Model Utility Scores [%] $\uparrow$ Attack Success Rate [%] $\downarrow$
@@
Zephyr-7B No Defense (HF) 81.0 55.2 56.2 100.0 61.4 98.8 85.0 87.5 85.0 97.5 90.0 85.0 100.0 100.0
CAT [7] (HF) 78.2 51.1 54.8 97.5 55.4 50.8 2.5 40.0 42.5 42.5 2.5 5.0 5.0 70.0
PAP-AT 82.3 54.2 56.4 97.5 54.1 94.0 17.5 2.5 5.0 15.0 2.5 55.0 57.5 77.5
DualAT 81.8 54.4 56.1 85.0 54.2 47.2 2.5 2.5 10.0 15.0 0.0 10.0 2.5 22.5
MixAT 81.4 54.0 55.8 97.5 54.3 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
MixAT Static 82.3 55.1 56.0 95.0 56.3 73.6 2.5 7.5 5.0 20.0 2.5 10.0 7.5 25.0
Llama3-8B No Defense (HF) 79.1 49.1 60.8 100.0 73.3 98.0 25.0 45.0 47.5 67.5 22.5 47.5 82.5 90.0
CAT [7] (R) 79.7 50.9 58.0 65.0 65.7 48.4 0.0 30.0 47.5 70.0 0.0 7.5 5.0 82.5
PAP-AT 81.1 51.9 60.2 100.0 58.6 84.4 22.5 2.5 15.0 22.5 10.0 52.5 40.0 70.0
DualAT 80.7 50.6 59.9 67.5 57.1 32.8 0.0 10.0 7.5 25.0 0.0 20.0 0.0 37.5
MixAT 80.4 50.1 59.1 85.0 55.6 40.0 0.0 0.0 2.5 2.5 0.0 22.5 0.0 25.0
MixAT Static 81.2 52.5 60.0 90.0 61.0 56.0 12.5 12.5 22.5 30.0 5.0 40.0 25.0 55.0
@@ B.9 Evaluating MixAT against other attacks
MixAT is evaluated against other attacks, including BEAST [37] and I-FSJ [38]. The results are in Tables 14 and 14, respectively. We observe that MixAT is robust against these attacks as well. This demonstrates that even though MixAT is trained using a combination of PAP and continuous attacks, the robustness properties generalize to a wide range of attacks.
@@
B.10 Examining malicious requests

### v2

@@ B.2 Discussion on Randomness
Randomness of attack generation All of the adversarial attacks we use in our experiments have a random component in their generation and evaluation process, since including random sampling from a distribution (GCG) or querying LLMs with a non-zero temperature (TAP, PAIR, AutoDAN). The adversarial attacks are evaluated by LLM-as-a-judge [34], which might cause inaccuracies. Randomness can yield different results when evaluating the robustness of the same model. To measure the extent of this randomness, we generate the adversarial samples for a single model multiple times and evaluate the robustness of the model on each set of samples. The results in Table˜4 show that the variance in the ASR scores is relatively low in the case of Zephyr-7B Base and MixAT model, while CAT results (PAP, TAP, PAIR) have significant randomness. This can stem from that the evaluation results of CAT indicating moderate robustness, which can only defend against weaker adversarial attacks. Therefore, this huge variance might indicate one limitation of LLM adversarial attacks, namely that it is difficult to generate similarly strong adversarial examples.
@@
R1 64.4 97.5 53.8 88.4 2.5 10.0 12.5
R2 63.8 97.5 52.2 74.0 0.0 12.5 12.5
R3 64.2 95.0 51.7 62.8 7.5 7.5 12.5
Avg 64.1 96.7 52.6 75.1 3.3 10.0 12.5
Std 0.31 1.44 1.1 12.83 3.82 2.50 0.00
Randomness of training We also examine the effect of randomness in the training process. We train the same model with different random seeds and evaluate the robustness of the models against adversarial attacks. The results in Table˜5 show that the variance induced by the randomness in the MixAT training procedure of Zephyr-7B has a relatively small magnitude, except the XSTest results. We would like to examine the causes of XSTest randomness in future work.
@@ B.3 Impact of Model Quantization
To examine the impact of model quantization, we train several models without 4-bit quantization and evaluate both the base and 4-bit quantized versions of them. The results (Table˜6) suggest, that models trained without 4-bit quantization are less robust than the models which were trained in 4-bit quantized form (higher ALO-ASR).
@@
Model
Quantization Utility Benchmarks [%] Attack Success Rate [%]
@@
Zephyr-7B
Base
Eval 81.0 55.2 56.2 100.0 60.3 98.8 85.0 87.5 85.0 97.5 90.0 85.0 100.0 100.0
None 81.3 57.3 58.1 100.0 61.8 98.8 77.5 72.5 90.0 100.0 95.0 90.0 100.0 100.0
Zephyr-7B
CAT
Train + Eval 78.2 50.5 54.5 95.0 52.7 50.0 0.0 25.0 27.5 55.0 0.0 12.5 0.0 67.5
Eval 77.7 50.4 55.0 90.0 53.5 47.6 0.0 45.0 55.0 65.0 0.0 12.5 0.0 87.5
None 79.2 51.5 56.9 90.0 54.4 50.8 0.0 52.5 52.5 75.0 0.0 25.0 2.5 90.0
Zephyr-7B
MixAT
Train + Eval 81.4 54.0 55.8 97.5 52.2 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
Eval 81.9 54.4 56.3 92.5 51.1 55.2 0.0 2.5 2.5 12.5 0.0 5.0 0.0 15.0
None 82.4 57.4 57.9 92.5 54.3 54.8 0.0 10.0 10.0 12.5 0.0 5.0 0.0 22.5
@@ B.4 Evaluating on more samples
In all our experiments we evaluate the models on a subset of 40 samples from the HarmBench test set. To assess the generalization of our models, we evaluate some of the Zephyr-7B models on the full test set of the HarmBench dataset. However, we still exclude the copyright samples as before, and we have a total of 240 samples. Since generating GCG attacks for all 240 samples would take 6 days on a single A100 GPU for each evaluated model, we only evaluate the models using the other attacks. The results in Table˜7 show that the models achieve similar robustness scores on the full test set as on the subset used in the main experiments. This indicates that evaluating on the 40-sample subset is a good proxy for the full HarmBench test set.
We also evaluate the robustness of the Zephyr models on a subset of 40 samples from the harmful prompts of the XSTest dataset. The results in Table˜8 show that the general robustness trends are similar across the two datasets, when considering the same attack.
@@
Model
Harmbench
Samples
Attack Success Rate [%]
@@ B.5 Transferability of attacks
Prior work [4] showed that adversarial attacks generated for non-adversarially-trained models transfer well to many other models. Here, we want to assess if adversarial attacks generated for adversarially-trained models exhibit the same trend. To do this, we generate adversarial samples using TAP, PAIR, and GCG attacks for different Zephyr-based models — undefended Z-Base, as well as, Z-CAT, and our Z-MixAT. We evaluate these sets of generated attacks on a large set of models both based on Zephyr and Llama3 and compare their efficiency versus Direct Requests and attacks specifically generated for the models. The results in Table˜9 show that almost all transferred attacks achieve worse performance compared to their targeted counterparts, with some transferred attacks being even less successful than the direct requests. Further, MixAT shows incredible resilience against all adversarial attacks, regardless of their origin.
@@
Defender
Model
Direct R.
ASR [%]
ASR [%] of TAP when target is ASR [%] of PAIR when target is ASR [%] of GCG when target is
@@ B.6 Utility benchmarks without chat template
Recent work highlights a failure mode in previous works, that evaluating LLMs on robustness and utility benchmark inconsistently with or without chat template might produce unrealistic results [7], since in real-world scenarios the settings of text generation do not depend on the content of prompts or whether they are malicious or ordinary requests. To enhance consistency, in this section we present ARC-E, ARC-C and MMLU results of LLama3 and Zephyr-7B with their default chat template Table˜10.
@@
Evaluated
Model
Chat Utility Benchmarks [%]
@@ B.7 MixAT on other models
To demonstrate the generalizability of our method, MixAT has been evaluated on additional models. We aimed to assess MixAT on an adversarially less robust (Mistral-7B[36]) and more robust (Llama3.1-8B [28]) base models. The evaluation results are in Table˜11. These are similar to our original findings based on Zephyr-7B, LLama3-8B and Qwen2.5 models. We observe that MixAT drastically increases the robustness of the base model in all of the cases, with a slight drop in utility.
@@
Model
Utility Scores [%] $\uparrow$ Attack Success Rate [%] $\downarrow$
@@
Mistral-7B
No Defense (HF) 79.7 49.9 52.5 100.0 57.9 99.2 80.0 77.5 90.0 95.0 95.0 80.0 100.0 100.0
CAT [7] (R) 79.8 50.0 52.8 90.0 53.6 68.8 2.5 60.0 80.0 80.0 7.5 37.5 32.5 95.0
PAP-AT 80.1 51.3 52.9 97.5 52.2 85.2 12.5 10.0 50.0 65.0 30.0 50.0 100.0 100.0
MixAT 79.8 50.3 52.8 92.5 48.8 55.2 2.5 7.5 22.5 25.0 0.0 40.0 10.0 52.5
Llama3.1-8B
No Defense (HF) 80.2 49.9 64.0 100.0 71.1 94.4 30.0 67.5 55.0 77.5 97.5 57.5 100.0 100.0
CAT [7] (R) 80.6 51.7 61.5 90.0 64.9 73.2 0.0 37.5 50.0 75.0 2.5 20.0 5.0 82.5
PAP-AT 81.3 51.1 64.1 100.0 61.4 92.4 15.0 5.0 12.5 22.5 7.5 55.0 35.0 67.5
MixAT 81.3 51.5 63.4 92.5 58.1 59.2 0.0 0.0 2.5 10.0 0.0 5.0 0.0 12.5
@@ B.8 MixAT Static training ablation results
We present the results of opur ablation using statically generated adversarial attacks in the training procedure of MixAT. The evaluation results are in Table˜12. We observe that while the static model has slightly higher utility, this comes at the cost of much worse robustness. This reaffirms the effectiveness of MixAT’s dynamic attack component.
@@
Model
Utility Scores [%] $\uparrow$ Attack Success Rate [%] $\downarrow$
@@
Zephyr-7B
No Defense (HF) 81.0 55.2 56.2 100.0 60.3 98.8 85.0 87.5 85.0 97.5 90.0 85.0 100.0 100.0
CAT [7] (HF) 78.2 51.1 54.8 97.5 52.8 50.8 2.5 40.0 42.5 42.5 2.5 5.0 5.0 70.0
PAP-AT 82.3 54.2 56.4 97.5 52.6 94.0 17.5 2.5 5.0 15.0 2.5 55.0 57.5 77.5
DualAT 81.8 54.4 56.1 85.0 51.1 47.2 2.5 2.5 10.0 15.0 0.0 10.0 2.5 22.5
MixAT 81.4 54.0 55.8 97.5 52.2 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
MixAT Static 82.3 55.1 56.0 95.0 54.0 73.6 2.5 7.5 5.0 20.0 2.5 10.0 7.5 25.0
Llama3-8B
No Defense (HF) 79.1 49.1 60.8 100.0 68.6 98.0 25.0 45.0 47.5 67.5 22.5 47.5 82.5 90.0
CAT [7] (R) 79.7 50.9 58.0 65.0 61.6 48.4 0.0 30.0 47.5 70.0 0.0 7.5 5.0 82.5
PAP-AT 81.1 51.9 60.2 100.0 57.5 84.4 22.5 2.5 15.0 22.5 10.0 52.5 40.0 70.0
DualAT 80.7 50.6 59.9 67.5 54.9 32.8 0.0 10.0 7.5 25.0 0.0 20.0 0.0 37.5
MixAT 80.4 50.1 59.1 85.0 53.1 40.0 0.0 0.0 2.5 2.5 0.0 22.5 0.0 25.0
MixAT Static 81.2 52.5 60.0 90.0 56.8 56.0 12.5 12.5 22.5 30.0 5.0 40.0 25.0 55.0
@@ B.9 Evaluating MixAT against other attacks
MixAT is evaluated against other attacks, including BEAST [40] and I-FSJ [41]. The results are in Tables˜14 and 14, respectively. We observe that MixAT is robust against these attacks as well. This demonstrates that even though MixAT is trained using a combination of PAP and continuous attacks, the robustness properties generalize to a wide range of attacks.
@@
Table 15: QueryAttack ASR on Qwen 2.5 14B
Base Model Method ASR $\downarrow$
Qwen 2.5 14B Base (HF) 0.965
CAT (R) 0.002
LAT KL (R) 0.021
MixAT 0.000
MixAT+GCG 0.000
Table 16: ArtPrompt ASR on Llama3-8B
Base Model Method ASR $\downarrow$
Llama3-8B Base (HF) 0.68
CAT (R) 0.02
LAT (HF) 0.02
MixAT 0.00
MixAT+GCG 0.00
B.10 MixAT + GCG with static GCG prompts
Since generating GCG attacks on-the-fly during training requires significant time, we further experimented with using static GCG examples created prior to the training for this purpose. Statistically generating all the GCG samples for the whole training set takes 33.5 hours on a single H200 GPU for Llama3-8B (36.6 hours for Zephyr), which makes this singular experiment significantly more expensive than our dynamic MixAT + GCG training. On the other hand, once we have generated the static GCG samples, we can run multiple training runs incorporating them. Our results in Table˜17 show that incorporating static GCG samples into MixAT training indeed shows similar benefits with the dynamically generated ones, providing an interesting avenue for future model training (datasets).
Table 17: Comparing MixAT + GCG with MixAT + static GCG on Zephyr and Llama3 models.
Model
Utility Scores [%] $\uparrow$ Attack Success Rate [%] $\downarrow$
ARCe ARCc MMLU Hless MTB XST D.R. PAP TAP PAIR A.DAN GCG H.Jail ALO
Zephyr-7B
CAT 78.2 51.1 54.8 97.5 52.8 50.8 2.5 40.0 42.5 42.5 2.5 5.0 5.0 70.0
MixAT (PAP+CAT) 81.4 54.0 55.8 97.5 52.2 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
MixAT + GCG 81.6 54.5 55.9 92.5 51.1 56.4 2.5 0.0 2.5 5.0 0.0 2.5 2.5 7.5
MixAT + static GCG 82.1 55.2 56.1 97.5 51.4 60.0 2.5 0.0 2.5 12.5 2.5 5.0 2.5 15.0
Llama3-8B
CAT 79.7 50.9 58.0 65.0 61.6 48.4 0.0 30.0 47.5 70.0 0.0 7.5 5.0 82.5
MixAT (PAP+CAT) 80.4 50.1 59.1 85.0 53.1 40.0 0.0 0.0 2.5 2.5 0.0 22.5 0.0 25.0
MixAT + GCG 80.1 48.7 58.5 92.5 50.7 47.6 0.0 0.0 5.0 7.5 0.0 2.5 0.0 15.0
MixAT + static GCG 79.1 48.4 56.9 87.5 53.1 40.0 0.0 0.0 0.0 2.5 0.0 2.5 0.0 5.0
B.11 MixAT with GCG as discrete attack
To assess how PAP contributes to the success of MixAT, we additionally train Zephyr and Llama3 MixAT models with statically generated GCG discrete attacks as seeds instead of PAP (using dynamically generated GCG samples would make the training too long and expensive). The results are in Table˜18. Our results demonstrate that using GCG samples as seeds for continuous attacks can improve the model’s robustness over the CAT baseline, but the lower diversity of GCG samples is detrimental to the generalization of robustness properties compared to rephrasing attacks such as PAP, TAP, and PAIR. In our MixAT + GCG ablation, we have shown that combining PAP, GCG, and Continuous attacks can further improve robustness, suggesting that the diversity of the attacks used for training is crucial for broad robustness. On the other hand, the comparison between MixAT and our baseline DualAT shows that the stronger mixed (discrete + continuous) attacks are better for training than simply combining the two training losses in a multi-objective manner.
Table 18: MixAT with static GCG discrete attacks instead of PAP on Zephyr and Llama3 models.
Model
Utility Scores [%] $\uparrow$ Attack Success Rate [%] $\downarrow$
ARCe ARCc MMLU Hless MTB XST D.R. PAP TAP PAIR A.DAN GCG H.Jail ALO
Zephyr
CAT 78.2 51.1 54.8 97.5 52.8 50.8 2.5 40.0 42.5 42.5 2.5 5.0 5.0 70.0
MixAT (PAP+CAT) 81.4 54.0 55.8 97.5 52.2 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
MixAT (GCG+CAT) 81.9 54.6 56.1 90.0 51.8 49.2 0.0 37.5 32.5 50.0 0.0 2.5 0.0 70.0
Llama3
CAT 79.7 50.9 58.0 65.0 61.6 48.4 0.0 30.0 47.5 70.0 0.0 7.5 5.0 82.5
MixAT (PAP+CAT) 80.4 50.1 59.1 85.0 53.1 40.0 0.0 0.0 2.5 2.5 0.0 22.5 0.0 25.0
MixAT (GCG+CAT) 80.8 50.3 59.4 77.5 55.9 43.2 0.0 35.0 42.5 55.0 0.0 5.0 2.5 67.5
B.12 Varying the hyper-parameters of the continuous PGD attack
To validate our hyperparameter choices, we show how scaling the $\epsilon$ (0.5x, 2x), the number of steps (0.5x, 2x), and the step-size (0.5x, 2.0x) of the continuous PGD attacks used in MixAT affects the robustness and performance of the trained models. Note that, for a fair comparison, we scaled the step-size with the perturbation budget and inversely with respect to the number of steps. The results are shown in Table˜19. Overall, we observe consistent results on the utility benchmarks, with only minor, to be expected, deviations from the original values. For robustness, this trend continues; however, we find that generally the chosen tradeoff of original MixAT performs on the top end of all tested configurations, with especially high-step low learning rate experiencing a stark drop in ALO robustness.
Table 19: Impact of varying MixAT hyperparameters on Zephyr-7B. We report utility scores (higher is better) and attack success rates (ASR, lower is better) across multiple hyperparameters.
MixAT Variant
Utility Scores [%] $\uparrow$ Attack Success Rate [%] $\downarrow$
ARC-E ARC-C MMLU Harmless MT-Bench XSTest Direct R PAP TAP PAIR AutoDAN GCG H Jail ALO
Original 81.4 54.0 55.8 97.5 52.2 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
0.5x $\epsilon$, 0.5x lr 82.0 55.2 56.4 97.5 51.5 63.2 2.5 0.0 2.5 2.5 0.0 12.5 2.5 12.5
2x $\epsilon$, 2x lr 81.9 53.8 55.5 92.5 49.7 81.6 0.0 0.0 5.0 5.0 0.0 2.5 0.0 12.5
2x steps, 0.5x lr 81.8 55.1 56.0 95.0 52.2 69.6 0.0 7.5 7.5 20.0 0.0 15.0 0.0 32.5
0.5x steps, 2x lr 81.4 54.5 56.1 97.5 51.6 56.8 0.0 2.5 10.0 10.0 0.0 10.0 2.5 20.0
B.13 Robustness against fundamentally different attack modalities
To show the generalization of MixAT defense to entirely different types of attacks (i.e., neither classical adversarial examples (GCG) nor paraphrasing/role-playing jailbreaks (TAP, PAP, HumanJailbreak)), we conducted two additional experiments - one using code attacks from the QueryAttack dataset [42] on Qwen 2.5 14B models; and one using ASCII art-based attacks from the ArtPrompt dataset [43] on Llama-3 8B models. We chose the model families per attack so that the attacks were particularly potent against the base models, allowing us to compare the defense methods meaningfully. The results are shown in Table˜16 and Table˜16, highlighting that all defense methods significantly improve robustness against these different attack modalities, with MixAT consistently achieving the best robustness.
B.14 Examining malicious requests